a) How many different ways can be letters in each of the following words be arranged?

i)YAHOO

ii)GOOGLE

b)How many of the arrangements in (a) will have the letters ‘O ́together?


1
Expert's answer
2021-05-28T09:58:47-0400

(a) (i) Number of ways to arrange 'Y' 'A' 'H' 'O' 'O'

Total letters in the word = 5

Since O comes two times

So total ways to arrange = "\dfrac{5!}{2!}=60"


(ii) Number of ways to arrange 'G' 'O' 'O' 'G' 'L' 'E'

Total letters in the word = 6

Since O and G comes two times

So, total ways to arrange = "\dfrac{6!}{2!\cdot 2!}=180"


(b) When two O's are together

so we have to arrange 'Y' 'A' 'H' 'OO'

Hence total ways to arrange = 4! = 24


and also we have to arrange 'G' 'OO' 'G' 'L' 'E'

Hence total ways to arrange = "\dfrac{5!}{2!}=60"